Given percent ionization (2.967%) and the pKa value (6.0165), find the inital concentration of [HA]i and the pH value.
Answer #1

Answer:-

This question is solved by using simple concept of ionisation of weak acid and expression of Ka and its value.

The answer is given in the image,

Аммол.:- percent conization 22-967% pka = 6•o165 Ка= 10-Рka io6-01 65 a = 9. 624 XJ6-7 НА + H2.0 Hgot + (ия), -x x x Ка - сноHow from euation CD 9.627X10-7 = 2.967x6*240): 12 CuAli -2.967X10-2 CHA) 9,627x10-7 = 8.80 3xlo-4 CHA), 2 0-97033 (HAJi CHAJC